OpenStudy (anonymous):

A train needs 12 seconds to go through a 50 yard tunnel. The same train going at the same speed needs 20 seconds to go through a 170 yard tunnel. What is the length of the train?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

6 yards

OpenStudy (asnaseer):

I believe you can solve this as follows: 1. call the length of the train L. 2. to go completely through a 50 yard tunnel it took 12s. this means it travelled 50+L yards in 12s. 3. to go completely through a 170 yard tunnel it took 20s. this means it travelled 120+L yards in 20s. since it's speed was the same in both case, we get the equation:\[\frac{50+L}{12}=\frac{120+L}{20}\]solve this for L.

OpenStudy (shane_b):

Makes sense to me...except I believe it should be 170+L on the right equation.

OpenStudy (asnaseer):

oops - well spotted @Shane_B - thx :)

OpenStudy (anonymous):

can u help me with 1 more question?

OpenStudy (shane_b):

@asnaseer: Np :)

OpenStudy (asnaseer):

@azntiger627 just post each question in the usual maner and someone will come along to help. If I'm free I'll help as well.

OpenStudy (anonymous):

I posted one, mine is the third one down